# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ 5.2.2 Methodology Analysis Among Teams
Most of the proposed approaches leverage both spatial and temporal learning to capture long-term dependencies within ultrasound sequences. Examples include the use of Mamba modules in FiMoNet and LSTM networks in RecuVol. ResNet and EfficientNet are employed as backbone architectures across several methods. All models predict 6-DoF transformations. While most methods estimate frame-to-frame transformations, FlowNet predicts transforma-tions between non-adjacent (interval) frames. Regarding loss functions, the primary objective across methods is to
@@
Pre-processing N/A Resize image to 50% width and height Normalising, downsampling by 1.25 (for half of the final ensemble models) Normalising to $[0,1]$ Cropping; scaling to $[-1,1]$ Fine-tuned Biomedical CLIP on the training dataset
@@
Data sampling N/A Randomly sampling scans of different lengths, ranging from 60 to 180 Sequences of 16 consecutive frames of the same scan N/A Randomly sample ultrasound sequence with $k+1$ frames N/A
@@
Loss MSE loss on transformed points coordinates $L_{1}$ loss, Pearson correlation loss MSE loss on transformation parameters MSE loss MMAE loss, Correlation loss and Triplet loss MSE loss, Consistency loss, Projection loss ($\alpha=1$, $\beta=0.69$, $\gamma=0.67$)
@@
minimise the discrepancy between the predicted and ground truth transformation parameters, commonly using $L_{1}$, MSE, or Pearson correlation-based losses (e.g., FiMoNet, RecuVol). Additional loss formulations are also utilised: MoGLo-Net employs a triplet loss; FlowNet incorporates a point-based loss on transformed coordinates and an MSE loss between original and warped ultrasound images; and PLPPI integrates embedding consistency loss and projection loss on landmarks. Pre-training is adopted by three approaches: FiMoNet and RecuVol use ImageNet-pretrained weights, while PLPPI employs foundation model Biomedical CLIP. Ensemble learning is another common strategy. FiMoNet combines two distinct models; RecuVol aggregates eight models derived from two rounds of 5-fold cross-validation; and FlowNet selects three models from different training epochs.
Illustrated in Table LABEL:imp, all methods are trained in an end-to-end manner and utilise offline inference. Most teams employs the Adam optimizer, with two opting for AdamW. A uniform base learning rate of 1e-4 is used across all submissions, though learning rate scheduling varies, including approaches such as StepLR, ReduceLROnPlateau, and cosine annealing with warmup. Training epochs span from under 100 to 13,400. Batch size varies between 1 and 32. The teams utilise a variety of GPU configurations for model training, including single-GPU setups with NVIDIA Quadro GV100, RTX 3090, 4090, and A6000, as well as a dual-GPU setup with A40s. Training times range from 1.2 to 9.7 GPU days, depending on resources and setup. None of the teams report the use of external data during training. Standard preprocessing steps, including scaling, cropping, and normalisation, along with data augmentation techniques such as temporal sampling and flipping, are consistently applied.
@@ 5.3.1 Overall Performance
Tables 5 and 6 present the performance of each team, assessed using four evaluation metrics along with their corresponding normalised scores. Figs. 6(a) and 7 provide a graphical representation of the same data. The abbreviations FS, GS, LS, PS, and LMS refer to the final score, global score, local score, pixel score, and landmark score, respectively. The evaluation results demonstrate that composite metrics effectively capture the strengths and limitations of participating methods across multiple spatial levels. The top-performing methods are characterised by strong local accuracy, low landmark error, and balanced global performance, indicating the importance of combining spatial-temporal modeling and ensemble learning.
@@
FiMoNet leads in 3 out of 4 normalised scores, particularly in local scores (LS: $0.951\pm 0.074$) and frame-to-frame accuracy (LPE: $0.097\pm 0.014$, LLE: $0.084\pm 0.019$), reflecting the advantage of its use of Mamba for temporal modeling, Pearson correlation-based loss, and dual-model ensembling. Its relatively low runtime also highlights a favorable balance between accuracy and efficiency. Close behind, RecuVol shows more balanced performance across global and pixel-level metrics (e.g., PS: $0.835\pm 0.131$, GLE: $5.978\pm 3.719$), but slightly lower local precision than FiMoNet. This indicates that while its LSTM-based temporal modeling and extensive ensemble setup (eight models) improve robustness, it may not capture local spatial structures as effectively. FlowNet, although achieving the lowest global errors (GPE: $5.970\pm 3.523$, GLE: $5.167\pm 3.682$), ranks lower in local-related metrics. This suggests that its interval-based frame prediction strategy and point/image-based loss functions capture coarse alignment well, but are less suited for precise local alignment. Its high inference time also presents a practical limitation.
MoGLo-Net and PLPPI show lower performance across all scores, with notably low landmark ($0.551\pm 0.270$ and $0.322\pm 0.240$) and global scores ($0.548\pm 0.322$ and $0.272\pm 0.302$). This suggests that their strategies, such as triplet loss in MoGLo-Net and embedding/projection losses in PLPPI, may not compensate for the lack of strong temporal modeling or ensemble learning. The Baseline model shows comparatively lower performance across all metrics, particularly in local alignment (LS: $0.056\pm 0.106$) and landmark localisation (LLE: $0.118\pm 0.031$).
@@
The final score, which normalises performance based on global and local transfromations, on all pixel and landmark level errors, ranks FiMoNet ($0.852$) and RecuVol ($0.817$) highest, indicating superior overall accuracy. These methods employ temporal modeling (Mamba and LSTM, respectively), diverse loss functions, and ensemble strategies, suggesting that integrating spatial-temporal features with strong supervision contributes to consistent performance across all spatial scales. Disaggregated metrics reveal further insights. The global score, based on GPE and GLE, highlights models that excel in aligning entire ultrasound scan. FlowNet, despite ranking third overall, achieves the best GPE ($5.970$ mm), reflecting strong global transformation learning. However, its local score is substantially lower ($0.622$), indicating that precise local alignment is not guaranteed by low global error alone. In contrast, FiMoNet achieves the highest local score (LS: 0.951), suggesting its feature extraction strategy, fine-grained feature extraction at multiple scales, is particularly effective at capturing anatomical detail. The pixel-wise score and landmark score aggregate global-local accuracy at pixel and landmark levels respectively. FiMoNet and RecuVol lead in both scores, reflecting their attention to both dense field alignment and landmark-specific accuracy.
@@
2 RecuVol $6.858\pm 3.526$ $5.978\pm 3.719$ $0.101\pm 0.016$ $0.088\pm 0.021$
3 FlowNet $\mathbf{5.970\pm 3.523}$ $\mathbf{5.167\pm 3.682}$ $0.111\pm 0.016$ $0.096\pm 0.022$
@@
  • For the five normalised scores, in scan level, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between methods PLPPI and baseline ($p$-value $=0.035$) in global score, methods FlowNet and MoGLo-Net ($p$-value $=0.033$) in local score. In subject level, for the five normalised scores, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between methods FiMoNet and RecuVol ($p$-value $=0.003$) in final score, FiMoNet and RecuVol ($p$-value $=0.021$) in global score, PLPPI and baseline ($p$-value $=0.358$) in global score, FlowNet and MoGLo-Net ($p$-value $=0.087$) in local score, FiMoNet and RecuVol ($p$-value $=0.021$) in landmark score.
  • For the four error metrics, in scan level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between PLPPI and the baseline method ($p$-value $=0.037$) in GPE metric, and FlowNet and MoGLo-Net ($p$-value $=0.011$) in LPE metric. In subject level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between methods FiMoNet and RecuVol ($p$-value $=0.008$) in GPE metric, methods PLPPI and baseline ($p$-value $=0.348$) in GPE metric, methods FiMoNet and RecuVol ($p$-value $=0.021$) in GLE metric, methods PLPPI and baseline ($p$-value $=0.081$) in GLE metric, and methods FlowNet and MoGLo-Net ($p$-value $=0.043$) in LPE metric.
Figure 6: (a) Box plots illustrating the distribution of performance scores across all test cases for each evaluated method (FiMoNet, RecuVol, FlowNet, MoGLO-Net, PLPPI, and Baseline), expressed in four normalised scores (FS, GS, LS, PS, LMS). The central line in each box represents the median score, with box edges indicating the first and third quartiles, and whiskers extending to 1.5 times the interquartile range (IQR). (b) Pairwise statistical comparison of final scores across methods in subject level, expressed as $p$-values from significance tests. Values close to zero indicate statistically significant differences in performance.
@@
Figure 8: (a) Ranking stability of participating teams using bootstrap sampling (2,000 bootstrap samples, each equal to test set size). The area of each bubble corresponds to the relative frequency of the respective ranks observed across bootstrap samples. The median rank for each algorithm is represented by a black cross. (b) Sampling distributions of the mean final score for each algorithm, approximated using the Central Limit Theorem. The normal curves show the probability density of the sample mean, highlighting spread and separation between algorithm performances.
To assess the robustness of algorithm rankings, we conducted a bootstrap analysis using 2,000 resampled test sets. Specifically, each bootstrap sample was created by resampling the test cases with replacement, keeping the sample size unchanged. Fig. 8(a) visualises the resulting rankings using a blob plot. The size of each bubble is proportional to the relative frequency of the corresponding ranks obtained across bootstrap samples. The median rank for each algorithm is denoted by a black cross. Notably, all algorithms exhibit perfect ranking consistency: each algorithm achieves the same rank in all 2,000 bootstrap samples, resulting in a single bubble per algorithm with 100% frequency. This suggests that performance differences among the algorithms are highly stable under resampling. Fig. 8(b) shows the estimated sampling distributions of the mean final score for each algorithm [55]. Each distribution is modeled as a Gaussian (normal) curve, where the center of the curve corresponds to the empirical mean of the final score for that algorithm, calculated across all test cases. The spread of the curve is determined by the standard error of the mean, computed as the sample standard deviation divided by the square root of the number of test cases. The tightness and separation of these curves reflect the consistency and distinguishability of algorithm performance.
@@ 5.3.2 Team-wise Performance Comparison Across Scan Patterns
Figs. 9 and 10 present the normalised scores and raw error metrics, respectively, for each team across various scan patterns. Local metrics (LPE, LLE) are lowest for straight line shape scans across all teams and highest for S shape scans. GPE and GLE (global metrics) are more consistent across shapes but still slightly higher for S shape scans, especially for methods like MoGLo-Net and PLPPI. This indicates that straight line shape scans are more tractable for the evaluated methods, and S shape scans introduce greater accumulated drift and lower global consistency due to their complex trajectories. Overall, linear trajectories such as those in straight line shape scans are handled more robustly by current methods, while more complex paths, especially S shape scans, lead to significant degradation in both global and local performance. These findings highlight the importance of evaluating reconstruction robustness under diverse motion patterns and reinforce the need for algorithms that generalise effectively across varied scanning conditions. Probe orientation also influences performance: all methods show higher LPE and LLE in parallel scans, while GPE and GLE are comparable across orientations over all methods but MoGLo-Net and PLPPI. This indicates that the parallel orientation may induce more accumulated drift and local misalignment, possibly due to less frame-to-frame overlap. These results highlight the need for methods that are not only scale-robust but also less sensitive to orientations, capable of maintaining accuracy under diverse scanning configurations. Across most methods, proximal-to-distal scans result in slightly higher metric errors compared to distal-to-proximal scans. In contrast, the influence of arm side (left vs. right) appears minimal, with comparable performance observed across both groups. Overall, these trends reveal the impact of scan shape, orientation, and direction on method robustness, and highlight the importance of developing reconstruction algorithms that generalise well across diverse scanning conditions. Performance scores (FS, GS, LS, PS, LMS) remain consistent across different scan patterns, as they are normalised metrics tend to be independent of absolute values.
5.3.3 Performance Analysis Across Scan patterns (Pooled Over All Teams)
@@
Given that scan length (SL) is a critical factor affecting reconstruction performance, primarily due to the cumulative nature inherent in freehand ultrasound reconstruction, this section presents a performance analysis across all methods with respect to both scan patterns and scan length. The observed relationship between performance and scan patterns is consistent with the findings reported in Section 5.3.2.
@@
Figs. 11 and 12 show the distribution of performance scores and error metrics across all methods for each scan, with scans ordered by increasing scan length. The trend suggests that, in general, all error metrics tend to increase with scan length, indicating a potential deterioration in both global trajectory accuracy and local alignment. However, this pattern is not strictly consistent across all cases, and some scans show minimal or no significant degradation. This overall tendency indicates that longer sequences may lead to greater accumulated drift, which affects performance at both global and local spatial scales. Performance scores (FS, GS, LS, PS, LMS) remain relatively stable across scan lengths, as they are normalised metrics designed to reduce sensitivity to overall error magnitude.
Fig. 13 presents the distribution of performance scores and error metrics across all methods for each subject, with subjects ordered by increasing scan length. The performance scores remain relatively stable across subjects. The error metrics demonstrate a positive correlation with scan length, increasing in subjects with longer scans.
Figs. 14 and 15 illustrate the distribution of performance scores and error metrics across individual scans, grouped by three scanning protocols: Straight line shape, C shape, and S shape. Within each protocol, scans are ordered by increasing scan length, enabling assessment of both protocol-specific and length-dependent trends. Across all three protocols, the normalised performance scores (FS, GS, LS, PS, LMS) generally remain within a consistent range, as expected. For the error metrics (GPE, GLE, LPE, LLE), global metrics (GPE, GLE) increase progressively with scan length in all scanning protocols. Local errors (LPE and LLE) show stable trend, suggesting that local alignment is comparatively less sensitive than global consistency. Comparing across scanning protocols, the straight line shape scans yield the lowest and most stable error values across most metrics. In contrast, the S shape scans are associated with the largest median scores and highest variability. C shape scans fall between these extremes, showing moderate errors and variability. Notably, for scans of comparable length, S shape sequences still perform worse, suggesting that both scan length and path complexity jointly negatively impact method effectiveness.
Figs. 16 and 17 present the distribution of performance scores and error metrics across all evaluated methods for scans categorised by probe orientation, parallel and perpendicular, and ordered by increasing scan length within each group. This configuration allows analysis of both orientation-dependent effects and scan length sensitivity. Performance scores are stable. Both global errors and local errors increase with scan length across both probe orientations. However, the increase is more noticeable in the parallel group, where both the magnitude and variability of errors are significantly higher. Overall, these findings demonstrate that both scan length and probe orientation significantly affect reconstruction performance. While performance degrades with increasing scan length in both settings, scans acquired in a parallel orientation exhibit higher global and local errors, along with greater performance variability.
Figs. 18 and 19 show the distribution of performance scores and error metrics across all methods for individual scans, categorised by the scanned arm (left or right) and arranged in ascending order of scan length. Across both groups, normalised scores (FS, GS, LS, PS, LMS) remain within a consistent range overall. Error metrics (GPE, GLE, LPE, LLE) consistently increase with scan length across both left and right arm scans. What is more, the error metrics are generally consistent between left and right arm scans, suggesting that the arm being scanned does not systematically bias metric magnitudes across methods compared to scan length.
@@
Figure 9: Normalised scores for each team across various scan patterns: straight-line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal and proximal-to-distal.
@@
Figure 10: Raw error metrics for each team across various scan patterns: straight-line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal and proximal-to-distal.
@@
Figure 12: Error metrics (GPE, GLE, LPE, LLE) for individual scans across all methods. All metrics are in millimeters (mm). Scans are sorted in ascending order of scan length (SL).
Figure 13: (a) Performance scores for individual subjects across all methods, arranged in ascending order of scan length. Higher scores indicate better performance. (b) Error metrics for individual subjects across all methods. Lower error indicates better performance. All metrics are in millimeters (mm). Scans are sorted in ascending order of scan length.
Figs. 20 and 21 explore the influence of scanning direction on performance, comparing scans acquired in a distal-to-proximal (DtP) versus proximal-to-distal (PtD) direction. Both global and local errors increase progressively with scan length in both directions, but with higher magnitudes and greater variability observed in the PtD group. In particular, GPE and GLE values are noticeably elevated in longer PtD scans, indicating more noticeable drift and loss of global alignment. Local metrics (LPE and LLE) also follow this trend but remain comparatively bounded, supporting the idea that global performance is more sensitive to scanning strategy and sequence length. These findings suggest that scanning direction influences the spatial continuity and reconstruction performance of sequential frames. The consistent increase in both global and local errors with scan length across both directions underscores the well-established challenge of long-sequence reconstruction. However, the amplified degradation in PtD scans indicates that the choice of scanning direction is a non-negligible factor in performance, especially for longer or more complex trajectories. Overall, these results highlight the importance of considering scanning direction as a variable in both evaluation and design of spatial tracking reconstruction systems. Future research may investigate direction-sensitive strategies for drift correction to improve robustness across scan patterns.

### v2

@@ 5.2.2 Methodology Analysis Among Teams
Most of the proposed approaches leverage both spatial and temporal learning to capture long-term dependencies within ultrasound sequences. Examples include the use of Mamba modules in FiMoNet and LSTM networks in RecuVol. ResNet and EfficientNet are selected as backbone architectures across several methods. All models predict 6-DoF transformations. While most methods estimate frame-to-frame transformations, FlowNet predicts transformations between non-adjacent (interval) frames. Regarding loss functions, the primary objective across methods is to minimise the discrepancy between the predicted and ground truth transformation parameters, commonly using $L1$, MSE, or Pearson correlation-based losses (e.g., FiMoNet, MoGLo-Net). Additional loss formulations are also utilised: MoGLo-Net uses a triplet loss; FlowNet uses a point-based loss on transformed coordinates and an MSE loss between original and warped ultrasound images; and PLPPI integrates embedding consistency loss and MSE loss on landmark coordinates. Pre-training is adopted by three approaches: FiMoNet and RecuVol use ImageNet-pretrained weights, while PLPPI uses foundation model Biomedical CLIP. Ensemble learning is another common strategy. FiMoNet combines two distinct models; RecuVol aggregates eight models derived from two rounds of 4-fold cross-validation; and FlowNet selects three models from different training epochs and also combines predictions from sequences with different offsets.
@@
Pre-processing N/A Resize image to 50% width and height Normalising, downsampling by 1.25 (for half of the final ensemble models) Normalising to $[0,1]$ Cropping; scaling to [1,1][-1,1] Fine-tuned Biomedical CLIP on the training dataset
@@
Data sampling N/A Randomly sampling scans of different lengths, ranging from 60 to 180 Sequences of 16 consecutive frames of the same scan N/A Randomly sample ultrasound sequence with 5 frames N/A
@@
Loss MSE loss on transformed points coordinates $L1$ loss and Pearson correlation loss on transformation parameters MSE loss on transformation parameters MSE loss on transformed point coordinates, transform parameters and warped images MMAE loss, Correlation loss and Triplet loss MSE loss, Consistency loss, Projection loss ($\alpha=1$, $\beta=0.69$, $\gamma=0.67$)
@@
Illustrated in Table LABEL:imp, all methods are trained in an end-to-end manner and utilise offline inference. Most teams adtopt the Adam optimizer, with two opting for AdamW. A uniform base learning rate of 1e-4 is used across all submissions, though learning rate scheduling varies, including approaches such as StepLR, ReduceLROnPlateau, and cosine annealing with warmup. Training epochs span from under 100 to 13,400. Batch size varies between 1 and 32. The teams utilise a variety of GPU configurations for model training, including single-GPU setups with NVIDIA Quadro GV100, RTX 3090, 4090, and A6000, as well as a dual-GPU setup with A40s. Training times range from 1.2 to 9.7 GPU days, depending on resources and setup. None of the teams report the use of external data during training. Standard preprocessing steps, including scaling, cropping, and normalisation, along with data augmentation techniques such as temporal sampling and flipping, are commonly applied.
@@ 5.3.1 Overall Performance
Tables 5 and 6 present the performance of each team, assessed using four evaluation metrics along with their corresponding normalised scores, as defined in Section 4.3 and 4.4. Figs. 7(a) and 8 provide a graphical representation with more detailed distributions of the evaluation metrics and scores. The abbreviations FS, GS, LS, PS, and LMS refer to the final score, global score, local score, pixel score, and landmark score, respectively. The evaluation results demonstrate that composite metrics effectively capture the strengths and limitations of participating methods across multiple spatial levels.
@@
FiMoNet leads in 3 out of 4 normalised scores, particularly in local scores (LS: $0.951\pm 0.074$), and in two unnormalised metrics about frame-to-frame accuracy (LPE: $0.097\pm 0.014$, LLE: $0.084\pm 0.019$), reflecting the advantage of its use of Mamba for temporal modeling, Pearson correlation-based loss, and dual-model ensembling. Its relatively low runtime also highlights a favorable balance between accuracy and efficiency. Close behind, RecuVol achieves better global metrics (GPE: $6.858\pm 3.526$, GLE: $5.978\pm 3.719$), but slightly lower local precision than FiMoNet. This indicates that while its LSTM-based temporal modeling and extensive ensemble setup (eight models) improve robustness, it may not capture local spatial structures as effectively. FlowNet, although achieving the lowest global errors (GPE: $5.970\pm 3.523$, GLE: $5.167\pm 3.682$), ranks lower in local-related metrics. This suggests that its interval-based frame prediction strategy and point/image-based loss functions capture coarse alignment well, but are less suited for precise local alignment. Its high inference time also presents a practical limitation. MoGLo-Net and PLPPI show lower performance across all scores, with notably low landmark ($0.551\pm 0.270$ and $0.322\pm 0.240$) and global scores ($0.548\pm 0.322$ and $0.272\pm 0.302$). This suggests that their strategies, such as triplet loss in MoGLo-Net and embedding/projection losses in PLPPI, may not compensate for the lack of strong temporal modeling or ensemble learning. The Baseline model shows the lowest performance across all metrics, particularly in local alignment (LS: $0.056\pm 0.106$), but requires the shortest run time due to its simplicity.
@@
The final score, which normalises performance based on global and local transformations, on all pixel and landmark level errors, ranks FiMoNet ($0.852$) and RecuVol ($0.817$) highest, indicating superior overall accuracy. These methods employ temporal modeling (Mamba and LSTM, respectively) and ensemble strategies, suggesting that integrating spatial-temporal features and model ensembling contributes to consistent performance across spatial scales. Disaggregated metrics reveal further insights. The global score, based on GPE and GLE, highlights models that excel in aligning entire ultrasound scan. FlowNet, despite ranking third overall, achieves the best GPE ($5.970$ mm), reflecting strong global transformation learning. However, its local score is substantially lower ($0.622$), indicating that precise local alignment is not guaranteed by low global error alone. In contrast, FiMoNet achieves the highest local score (LS: 0.951), suggesting its fine-grained feature extraction at multiple scales is particularly effective at capturing anatomical detail.
@@
2 RecuVol $6.858\pm 3.526$ $5.978\pm 3.719$ $101\pm 0.016 0.0$$088\pm 0.021$
3 FlowNet $\mathbf{5.970\pm 3.523}$ $\mathbf{5.167\pm 3.682}$ $111\pm 0.016 0.0$$096\pm 0.022$
@@
  • For the five normalised scores, at scan level, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between PLPPI and baseline ($p$-value =0.035=0.035) in global score, and FlowNet vs MoGLo-Net ($p$-value =0.033=0.033) in local score. At subject level, for the five normalised scores, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between FiMoNet and RecuVol ($p$-value =0.003=0.003) in final score, FiMoNet and RecuVol ($p$-value =0.021=0.021) in global score, PLPPI and baseline ($p$-value =0.358=0.358) in global score, FlowNet and MoGLo-Net ($p$-value =0.087=0.087) in local score, FiMoNet and RecuVol ($p$-value =0.021=0.021) in landmark score.
  • For the four error metrics, at scan level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between PLPPI and the baseline method ($p$-value =0.037=0.037) in GPE metric, and FlowNet vs MoGLo-Net ($p$-value =0.011=0.011) in LPE metric. At subject level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between FiMoNet and RecuVol ($p$-value =0.008=0.008) in GPE metric, PLPPI and baseline ($p$-value =0.348=0.348) in GPE metric, FiMoNet and RecuVol ($p$-value =0.021=0.021) in GLE metric, PLPPI and baseline ($p$-value =0.081=0.081) in GLE metric, and FlowNet vs MoGLo-Net ($p$-value =0.043=0.043) in LPE metric.
Figure 7: (a) Box plots illustrating the distribution of performance scores (FS, GS, LS, PS, LMS) across all test cases for each evaluated method (FiMoNet, RecuVol, FlowNet, MoGLO-Net, PLPPI, and Baseline). The central line in each box represents the median score, with box edges indicating the first and third quartiles, and whiskers extending to 1.5 times the interquartile range (IQR). (b) $p$-values for pairwise, subject-level statistical comparison of final scores across methods. Values close to zero indicate statistically significant differences in performance.
@@
Figure 9: (a) Ranking stability of participating teams using bootstrap sampling (2,000 bootstrap samples, each equal to test set size). The area of each bubble corresponds to the relative frequency of the respective ranks observed across bootstrap samples. The median rank for each algorithm is represented by a black cross. (b) Sampling distributions of the mean final score for each algorithm, approximated using the Central Limit Theorem. The curves show the probability density of the sample mean, highlighting spread and separation between algorithm performances.
To assess the robustness of algorithm rankings, we conducted a bootstrap analysis using 2,000 resampled test sets. Specifically, each bootstrap sample was created by resampling the test cases with replacement, keeping the sample size unchanged. Fig. 9(a) visualises the resulting rankings using a blob plot. The size of each bubble is proportional to the relative frequency of the corresponding ranks obtained across bootstrap samples. The median rank for each algorithm is denoted by a black cross. Notably, all algorithms exhibit perfect ranking consistency: each algorithm achieves the same rank in all 2,000 bootstrap samples, resulting in a single bubble per algorithm with 100% frequency. This suggests that performance differences among the algorithms are highly stable under resampling. Fig. 9(b) shows the estimated sampling distributions of the mean final score for each algorithm [97]. Each distribution is modeled as a Gaussian (normal) distribution, where the center of the curve corresponds to the empirical mean of the final score for that algorithm, calculated across all bootstrap samples. The standard deviation is determined by the standard error of the mean, computed as the sample standard deviation divided by the square root of the number of bootstrap samples. The tightness and separation of these curves provide insight into the consistency and distinguishability of algorithm performance rankings.
@@ 5.3.2 Team-wise Performance Comparison Across Scan Patterns
Figs. 10 and 11 present the normalised scores and raw error metrics, respectively, for each team across various scan patterns. Consistently for all teams, Straight line shape scans have the lowest local metrics (LPE, LLE) while S shape scans have the highest local metrics. For global metrics (GPE, GLE), different scan shapes show more consistent performance, although S shape scans still show slightly higher global errors, especially for methods like MoGLo-Net and PLPPI. This indicates that straight line shape scans are more tractable for the evaluated methods, while S shape scans are prone to greater drift and lower global consistency due to their complex trajectories. Overall, linear trajectories such as those in straight line shape scans are handled more robustly by current methods, while more complex paths, especially S shape scans, lead to significant degradation in both global and local performance. These findings highlight the importance of evaluating reconstruction robustness under diverse motion patterns and reinforce the need for algorithms that generalise effectively across varied scanning conditions.
Probe orientation also influences performance: all methods show higher LPE and LLE in parallel scans than in perpendicular scans, while GPE and GLE are comparable across orientations for all methods but MoGLo-Net and PLPPI. This indicates that the parallel orientation may induce more local misalignment, possibly due to less frame-to-frame overlap.
For most methods, proximal-to-distal scans result in slightly higher error metrics compared to distal-to-proximal scans. In contrast, the influence of arm side (left vs. right) appears minimal, with comparable performance observed across both groups. Overall, these trends reveal the impact of scan shape, orientation, and direction on method robustness, and highlight the importance of developing reconstruction algorithms that generalise well across diverse scanning conditions. As shown in Fig. 10, performance scores (FS, GS, LS, PS, LMS) remain consistent across different scan patterns, as they are normalised metrics tend to be independent of absolute values.
5.3.3 Aggregated Performance over All Teams and Influence of Scan Length
@@
Scan length (SL) is a critical factor affecting reconstruction performance. This section analyses the average performance of all methods with respect to both scan patterns and scan length. Overall, the observed relationship between performance and scan patterns is consistent with the findings reported in Section 5.3.2.
@@
Figs. 12 and 13 show the distribution of performance scores and error metrics across all methods for each scan, respectively. In the horizontal axis, the scans are sorted by increasing scan length. The trend suggests that, in general, all error metrics tend to increase with scan length, indicating a potential deterioration in both global trajectory accuracy and local alignment. However, this pattern is not strictly consistent across all cases, and some scans show minimal or no significant degradation. This overall tendency indicates that longer sequences may lead to greater drift, which affects performance at both global and local spatial scales. Performance scores (FS, GS, LS, PS, LMS) remain relatively stable across scan lengths, as they are normalised metrics designed to reduce sensitivity to overall error magnitude.
@@
Figure 10: Normalised scores for each team across various scan patterns: Straight line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal (DtP) and proximal-to-distal (PtD).
@@
Figure 11: Raw error metrics for each team across various scan patterns: Straight line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal (DtP) and proximal-to-distal (PtD).
Fig. 14 presents the distribution of performance scores and error metrics across all methods for each subject, with subjects sorted by increasing scan length. The performance scores remain relatively stable across subjects. The error metrics demonstrate a positive correlation with scan length, increasing in subjects with longer scans.
Figs. 15 and 16 illustrate the distribution of performance scores and error metrics across individual scans, grouped by three scanning protocols: Straight line shape, C shape, and S shape. Within each protocol, scans are ordered by increasing scan length, enabling assessment of both protocol-specific and length-dependent trends. Across all three protocols, the normalised performance scores (FS, GS, LS, PS, LMS) generally remain within a consistent range, as expected. For the error metrics, global metrics (GPE, GLE) increase progressively with scan length in all scanning protocols.
@@
Local errors (LPE and LLE) show stable trend, suggesting that local alignment is comparatively less sensitive than global consistency. Comparing across scanning protocols, the straight line shape scans yield the lowest and most stable error values across most metrics. In contrast, the S shape scans are associated with the largest error and highest variability. C shape scans fall between these extremes, showing moderate errors and variability. Notably, for scans of comparable length, S shape sequences still perform worse, suggesting again that both scan length and path complexity jointly negatively impact method effectiveness.
Figs. 17 and 18 present the distribution of performance scores and error metrics across all evaluated methods for scans categorised by probe orientation, parallel and perpendicular, respectively. Scans are sorted by their length within each group. This configuration allows analysis of both orientation-dependent effects and scan length sensitivity. Normalised performance scores are stable. Both global errors and local errors increase with scan length across both probe orientations. However, the increase is more noticeable in the parallel group, where both the magnitude and variability of errors are significantly higher. They demonstrate that both scan length and probe orientation significantly affect reconstruction performance.
Figs. 19 and 20 show the distribution of performance scores and error metrics across all methods for individual scans, categorised by the scanned arm (left or right) and arranged in ascending order of scan length. Across both groups, normalised scores remain within a consistent range overall. Error metrics consistently increase with scan length across both left and right arm scans. In addition, the error metrics are generally consistent between left and right arm scans, suggesting that the arm being scanned does not systematically influence the metric values across methods compared to scan length.
Figure 13: Error metrics (GPE, GLE, LPE, LLE) for individual scans across all methods. All metrics are in millimeters ($m$). Scans are sorted in ascending order of scan length (SL).
Figure 14: (a) Performance scores for individual subjects across all methods, arranged in ascending order of scan length. Higher scores indicate better performance. (b) Error metrics for individual subjects across all methods. Lower error indicates better performance. All metrics are in millimeters ($m$). Scans are sorted in ascending order of scan length.
Figs. 21 and 22 explore the influence of scanning direction on performance, comparing scans acquired in a distal-to-proximal (DtP) versus proximal-to-distal (PtD) direction. Both global and local errors increase progressively with scan length in both directions, but with higher magnitudes and greater variability observed in the PtD group. In particular, GPE and GLE values are noticeably elevated in longer PtD scans, indicating more noticeable drift and loss of global alignment. Local metrics (LPE and LLE) also follow this trend but remain comparatively bounded, supporting the idea that global performance is more sensitive to scanning strategy and sequence length. These findings suggest that scanning direction influences the spatial alignment and reconstruction performance of sequential frames. The amplified degradation in PtD scans indicates that the choice of scanning direction is a non-negligible factor in performance, especially for longer or more complex trajectories. Overall, these results highlight the importance of considering scanning direction as a variable in both evaluation and design of spatial tracking reconstruction systems. Future research may investigate direction-sensitive strategies for drift correction to improve robustness across scan patterns.